SPECTRAL ANALYSIS OF THE BENDING LEVELS OF ACETYLENE CATION

YUNG-CHING CHOU and \underline{YEN-CHU HSU}, INSTITUTE OF ATOMIC AND MOLECULAR SCI-ENCES, ACADEMIA SINICA, P.O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C..

Twenty-seven vibronic levels (v_{4}=0-6, v_{5}=0-2, K=0-3) of the ground electronic state of acetylene cation have been previously studied by 1+1’ two-color pulsed-electric field induced zero-kinetic energy (PFI-ZEKE) photoelectron spectroscopy^{a}. These levels, rotationally resolved, are about 0-3520 cm^{−1} above the zero-point energy level. Previously, only the Renner-Teller effects of the trans- and cis-bending vibration and the vibrational anharmonicity were included in the effective Hamiltonian^{b}; one standard deviation of the fit was 4.1 cm^{−1}, about six times of the measurement uncertainties. Large discrepancies between the observed and calculated frequencies were identified at combination levels. In this fit, effects such as vibrational ℓ−type doubling and second-order anharmonic interaction were also considered due to the near degeneracy of the trans- and cis-bending vibration. The results of the fit are satisfactory; most of the vibronic bands can be fitted within one standard deviation of 1.0 cm^{−1}, except the lower component of (v_{4}=0, v_{5}=1, K=0). The results of the fit will be presented and few observed vibronic levels, yet unassigned, will be discussed.

aShuenn-Jiun Tang,Ph. D. thesis, Chemistry Department, National Taiwan University, Taiwan, R. O. C., 1999.

bJ. M. Brown,J. Mol. Spectrosc.\underline{68},412(1977).